OpenStudy (anonymous):

the radii of two circles are in the ratio of 3 to 1.find the area of the smaller circle if the area of the larger circle is 27pi sq. inches

OpenStudy (turingtest):

do you know the formula for the area of a circle?

OpenStudy (anonymous):

area larger circle =pi*R^2=pi*27.....R^2=27......R=3sqrt3 smaller circle =.................................................r=R/3= 3sqrt3/3= sqrt3 area smaller circle =pi *r^2= pi *3